A circular oil spill has a radius of 2 miles. After a day, the radius of the oil spill increases by 3 miles. By how many square miles does the area of the oil spill increase? Round to the nearest thousand, if necessary


1
Expert's answer
2022-03-14T04:37:15-0400

"A_1=\pi2^2=4\pi sq.mi."

"A_2=\pi(2+3)^2=25\pi sq.mi."

"A_2-A_1=25\pi-4\pi=21\pi\approx 65.973 sq.mi."